# High Quality QCD Axion in the Standard Model

### v1

@@ II Anomaly-Free Discrete Gauge Symmetries
Since all the required operators contain only even powers of $\Phi$, this means that $\Phi$ naturally possesses a $\mathbb{Z}_{2}$ gauge symmetry. Only the field $\Phi$ is odd under the $\mathbb{Z}_{2}$ while all other fields transform as the $\mathbb{Z}_{2}$-even. This symmetry is anomaly-free, since it does not act on the chiral fermions. Thus, we impose this exact discrete $Z_{2}$ symmetry throughout this letter. We will show that this minimal extension of the SM naturally processes the PQ symmetry, with the field $\Phi$ playing the role of the PQ field, and elegantly resolves the axion high-quality problem.
@@ III High-Quality Axion
A higher dimensional operator involving $\Phi$ that couples $H_{1}$ and $H_{2}$ is allowed under the discrete symmetris,
@@
The QCD instantons generate an effective axion potential as $V_{a}=m_{a}^{2}F_{a}^{2}[1-\cos(\bar{\theta}+N_{D}a/F_{a})]$, which gives rise to the axion mass $m_{a}$. This dynamical axion field drives its potential to the minimum, that is, $N_{D}a/F_{a}=-\bar{\theta}$, thereby solving the strong CP problem. The axion mass is $m_{a}=[\sqrt{m_{u}m_{d}}/(mu+md)]\times N_{D}m_{\pi}f_{\pi}/F_{a}=N_{D}\times 5.7\,\mu\text{eV}\times(10^{12}\,\text{GeV}/F_{a})$, where up-quark mass $m_{u}=2.16\,$MeV, down quark mass $m_{d}=4.67\,$MeV, pion mass $m_{\pi}\simeq 140\,\mathrm{MeV}$, and pion decay constant $f_{\pi}=92\,\mathrm{MeV}$. Additionally, $N_{D}$ denotes the domain-wall number, which counts the number of distinct vacua and appears in the axion–gluon interaction as $\mathcal{L}=N_{D}aG\tilde{G}/F_{a}$ Kim:1986ax ; Kim:2008hd .
@@
However, in our setup, with the charge assignment (1,1)(-1,-1) for the $\Phi$, the lowest-dimensional leading term of the potential consistent with both $\mathbb{Z}_{3}$ and $\mathbb{Z}_{4}$ discrete symmetries is [^2] We have also allowed higher dimensional operator $(H_{1}^{\dagger}H_{2})^{2}\Phi^{4}$. However, the contribution to the extra mass for axion can be negligible because of the small vevs of $H_{1,2}$.,
@@ IV Two-Component Dark Matter
$m_{\chi}=\frac{1}{4!}\frac{\braket{\Phi}^{4}}{M_{\text{pl}}^{4}}=50\,\text{eV}\times\left(\frac{\braket{\Phi}}{2\times 10^{12}\,\text{GeV}}\right)^{4}.$ (12)
@@
shown as the red solid line in Fig. 1. Different from other benchmark models, our model has a well-defined prediction. In most of these predicted parameter space, axion is still the dominant DM component and can be tested by direct detection. In this mass range, haloscope experiments CAPP:2024dtx ; QUAX:2024fut ; Quiskamp:2024oet ; ADMX:2024xbv targeting for the axion-photon coupling have already begun to approach the KSVZ model parameter space, and will soon be able to probe the couplings predicted by our model, which are similar to those in DFSZ model [^3] The axion-photon coupling is determined by the ratio of the electromagnetic anomaly to the color anomaly coefficients. In our model, compared with the DFSZ model, although the PQ charge of fermions is doubled, this ratio remains the same.. Future haloscope experimental projections (grey dashed lines) could provide a definitive test of our scenario AxionLimits .

### v2

@@ II Anomaly-Free Discrete Gauge Symmetries
Since all the required operators contain only even powers of $\Phi$, this means that $\Phi$ naturally possesses a $\mathbb{Z}_{2}$ gauge symmetry. Only the field $\Phi$ is odd under the $\mathbb{Z}_{2}$ while all other fields transform as the $\mathbb{Z}_{2}$-even. This symmetry is anomaly-free, since it does not act on the chiral fermions. Thus, we impose this exact discrete $Z_{2}$ symmetry throughout this letter [^2] This $\mathbb{Z}_{2}$ is intact on the quality of the axion. It is assumed to guarantee the complete stability of the $\chi$ DM as explained later. . We will show that this minimal extension of the SM naturally processes the PQ symmetry, with the field $\Phi$ playing the role of the PQ field, and elegantly resolves the axion high-quality problem.
@@ III High-Quality Axion
A higher dimensional operator involving $\Phi$ that couples $H_{1}$ and $H_{2}$ is allowed under the discrete symmetries and the $U(1)$ PQ symmetry,
@@
The QCD instantons generate an effective axion potential as $V_{a}=m_{a}^{2}F_{a}^{2}[1-\cos(\bar{\theta}+N_{D}a/F_{a})]$, which gives rise to the axion mass $m_{a}$. This dynamical axion field drives its potential to the minimum, that is, $N_{D}a/F_{a}=-\bar{\theta}$, thereby solving the strong CP problem. The axion mass is $m_{a}=[\sqrt{m_{u}m_{d}}/(m_{u}+m_{d})]\times N_{D}m_{\pi}f_{\pi}/F_{a}=N_{D}\times 5.7\,\mu\text{eV}\times(10^{12}\,\text{GeV}/F_{a})$, where up-quark mass $m_{u}=2.16\,$MeV, down quark mass $m_{d}=4.67\,$MeV, pion mass $m_{\pi}\simeq 140\,\mathrm{MeV}$, and pion decay constant $f_{\pi}=92\,\mathrm{MeV}$. Additionally, $N_{D}$ denotes the domain-wall number, which counts the number of distinct vacua and appears in the axion–gluon interaction as $\mathcal{L}=N_{D}aG\tilde{G}/F_{a}$ Kim:1986ax ; Kim:2008hd .
@@
However, in our setup, with the charge assignment (1,1)(-1,-1) for the $\Phi$, the lowest-dimensional leading term of the potential consistent with both $\mathbb{Z}_{3}$ and $\mathbb{Z}_{4}$ discrete symmetries is [^3] We have also allowed higher dimensional operator $(H_{1}^{\dagger}H_{2})^{2}\Phi^{4}$ that explicitly breaks $U(1)$ PQ symmetry. However, the contribution to the extra mass for axion can be negligible because of the small vevs of $H_{1,2}$.,
@@ IV Two-Component Dark Matter
$m_{\chi}=\frac{1}{4!}\frac{\braket{\Phi}^{4}}{M_{\text{pl}}^{3}}=50\,\text{eV}\times\left(\frac{\braket{\Phi}}{2\times 10^{12}\,\text{GeV}}\right)^{4}.$ (12)
@@
shown as the red solid line in Fig. 1. Different from other benchmark models, our model has a well-defined prediction. In most of these predicted parameter space, axion is still the dominant DM component and can be tested by direct detection. In this mass range, haloscope experiments PhysRevLett.59.839 ; Wuensch:1989sa ; PhysRevD.42.1297 ; Hagmann:1996qd ; McAllister:2017lkb ; Quiskamp:2022pks ; Quiskamp:2023ehr ; CAPP:2024dtx ; QUAX:2024fut ; Quiskamp:2024oet ; ADMX:2024xbv targeting for the axion-photon coupling have already begun to approach the KSVZ model parameter space, and will soon be able to probe the couplings predicted by our model, which are similar to those in DFSZ model [^4] The axion-photon coupling is determined by the ratio of the electromagnetic anomaly to the color anomaly coefficients. In our model, compared with the DFSZ model, although the PQ charge of fermions is doubled, this ratio remains the same.. Future haloscope experimental projections (grey dashed lines) could provide a definitive test of our scenario AxionLimits .
@@ V Conclusion and Discussions
However, if we consider the BLB-L gauge symmetry instead of the $\mathbb{Z}_{4}$, we have to introduce a new Higgs field to generate the masses of the right-handed neutrinos. In this extension of the SM, the masses of the right-handed neutrinos are independent of the $PQ$ breaking scale and they can be much larger than the $PQ$ breaking scale such as $M_{N_{i}}\sim 10^{15}$ GeV. See the Appendix for details of the BLB-L extension of the SM.
@@ Acknowledgements
Appendix A Appendix A – BLB-L Extension
The PQ field $\Phi$ couples to the right-handed neutrinos $\bar{N}_{i}$, and its vacuum expectation value generates their Majorana masses $M_{\bar{N}_{i}}$. Consequently, the masses of $\bar{N}_{i}$ are related to the PQ breaking scale $F_{a}$. This feature is an appealing aspect of the $\mathbb{Z}_{4}\times\mathbb{Z}_{3}$ extension of the SM Langacker:1986rj .
However, if we consider the $U(1)_{B-L}$ gauge instead of introducing the $\mathbb{Z}_{4}$ symmetry, a new Higgs field is naturally introduced to generate the right-handed neutrino masses. In this BLB-L extension, the masses of the right-handed neutrinos are no longer correlated with the PQ scale and may be significantly larger, e.g. $M_{N_{i}}\sim 10^{15}\,$GeV. In this section, we emphasize that the high quality of the QCD axion can still be maintained even within the BLB-L extension of the SM.
In the $U(1)_{B-L}$ extended SM, the $\mathbb{Z}_{12}$ discrete symmetry can be still anomaly free including the Dai-Freed anomaly Dai:1994kq ; Yonekura:2016wuc if we introduce the three left-handed fermions $\chi_{i}$, too. Now, we assume $SU(3)\times SU(2)\times U(1)_{Y}\times U(1)_{B-L}\times\mathbb{Z}_{12}$ gauge theory. The charges of all fields are shown in Tab. 3 and Tab. 4. As for the fermion sector the discrete symmetry is equivalent to the original $\mathbb{Z}_{3}$.
Fermions $T$ $\bar{F}$ $\bar{N}$ $\chi$
BLB-L 1 -3 5 0
$\mathbb{Z}_{12}$ 4 4 4 -4
Table 3: The charges of fermions under $U(1)_{B-L}$ and $\mathbb{Z}_{12}$ symmetries.
Higgs $H_{1}$ $H_{2}$ $\Phi$ $\Phi^{\prime}$
BLB-L 2 2 0 -10
$\mathbb{Z}_{12}$ -4 4 -1 4
Table 4: The charges of Higgs under $U(1)_{B-L}$ and $\mathbb{Z}_{12}$ symmetries.
The right-handed neutrinos $\bar{N}_{i}$ obtain Majorana masses $M_{N}$ from the vev of $\Phi^{\prime}$, since we have $\bar{N}_{i}\bar{N}_{i}\Phi^{\prime}$ coupling. On the other hand, the PQ symmetry is spontaneously broken down by the vev of $\Phi$ while Eq. (4) and $\chi\chi\Phi^{4}$ coupling in Eq. (6) are unchanged. The PQ symmetry is explicitly broken by Eq. (7) and hence we can keep the high quality of the axion as in the $\mathbb{Z}_{4}\times\mathbb{Z}_{3}$ model.